OpenStudy (anonymous):

In triangle ABC, what is the length of side b if B = 10°, C = 135°, and c = 45? 3.24 3.33 3.82 11.05

OpenStudy (dtan5457):

all your questions just require the laws of sine. what are you not getting exactly?

OpenStudy (anonymous):

all the different formulas

OpenStudy (anonymous):

and how to set them up

OpenStudy (dtan5457):

the laws of sine don't change lol

OpenStudy (dtan5457):

you were able to do them just fine on the last question

OpenStudy (anonymous):

I had repost this question

OpenStudy (anonymous):

Ima try to find it

OpenStudy (dtan5457):

45/sin135=side B/sin10

OpenStudy (anonymous):

\[\frac{ 45 }{ \sin 135 }=side B/\sin10\]

OpenStudy (dtan5457):

correct

OpenStudy (anonymous):

63.63961030678927719607599258943641353563523439196266329295058...

OpenStudy (anonymous):

and then Im probably puting it into the calculator wrong yet again

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=45%2Fsin%28135

OpenStudy (anonymous):

http://openstudy.com/study#/updates/55340c68e4b07417d4e0bd2c this is what he showed me

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (anonymous):

@wio

OpenStudy (anonymous):

@Nnesha

OpenStudy (anonymous):

maybe that will help

OpenStudy (anonymous):

and it does

Nnesha (nnesha):

try again

OpenStudy (anonymous):

cause now I am getting http://www.calculatorsoup.com/calculators/geometry-plane/triangle-law-of-sines.php

Nnesha (nnesha):

\(\color{blue}{\text{Originally Posted by}}\) @DaWizjr \[\frac{ 45 }{ \sin 135 }=side B/\sin10\] \(\color{blue}{\text{End of Quote}}\) equal what ?

OpenStudy (anonymous):

11.05

Nnesha (nnesha):

yes

OpenStudy (anonymous):

thank god for this calculator

OpenStudy (anonymous):

to late now ;/

OpenStudy (anonymous):

:) thanks

Nnesha (nnesha):

gO_Od job!

OpenStudy (anonymous):

guess you dont need medals